The recommended adult dose of Elixophyllin®, a drug used to treat asthma, is 6 mg/kg of body mass. Calculate the dose in milligrams for a 180-lb person.

(c) If an automobile is able to travel 230. mi on 10.7 gal of gasoline, what is the gas mileage in km/L?

(d) A pound of coffee beans yields 50 cups of coffee (4 cups = 1 qt). How many milliliters of coffee can be obtained from 8 g of coffee beans?
1
Expert's answer
2018-01-26T14:56:07-0500
180 lb = 81.65 kg
Recommended dose of Elixophyllin = 6·81.65 = 489.9 mg
(c) 230 mi = 370.15 km
10.7 gal = 40.5 L
Gas mileage = 230 mi/10.7 gal = 370.15 km/40.5 L = 9.14 km/L
(d) 1 pound = 453.59 g
1 qt = 946.35 ml
50 cups = 12.5 qt = 11829.38 ml
453.59 g - 11829.38 ml
8 g - n (ml) of coffee
n (ml) of coffee = 8·11829.38/453.59 = 208.64 ml